Make New Entry,
Make Followup Entry
User name Markowitz
Log entry time 08:05:23 on July 05, 2010
Entry number 330176
keyword=Owl shift summary
Owl Shift Summary: July 4, 2010
- Pete Markowitz
- James Beacham
- Victor Maxwell
- Lawrence Selvy
- 00:00 Hall is in controlled access to remove the extra 100 ns
timing Bogdan put in the trigger to get accidentals. I *think* the 100
ns was added to T1 meaning the overall timing was not shifted, but that
anything put in coincidence with T1 would have to be an accidental.
Follow-up plot (entry 330112 by Alexandre) shows that we did get only
- 12:20 Hall is out of controlled access. Move to empty target and
ramp up septum. Beam spot on dump viewer is OK.
- 12:30 back to Pb-D1 target. Saw that T6 was left as OR rather than
AND. Ask for access to change. Septum off.
- 01:10 Septum back on, beam requested.
- 01:16 Start run 1713, Prescales 1e5 except PS6=1.
- 01:49 Start run 1714, Prescales 1e5 except PS4=1, PS6=1.
- 02:40 Ask MCC to do ion chamber calibration for upto 25 uA on Ta
- 04:00 Septum off for access. Put in sieve slits, narrow trigger
window. Putting in Ta target.
- 04:30 Septum on (out of hall, sieve in, trigger narrowed).
- 04:37 Start run 1717. Q1 trips. end run.
- 04:44 Start run 1718. Sieve in both arms, 1M events about half T1
and half T3.
- 04:52 Start run 1719. Same thing. Sergey adjusts rhrs lead glass
HV during run.
- 05:00 Start run 1720. Same thing with Sergey's new HV.
- 05:10 Ask for access to take out sieve. Septum off.
- 05:30 When beam comes back, we observe about the same rate as
before. Conclusion is that what we recorded was primarily pion
punchthroughs. Temporarily give up on optics at 2-pass.
- 05:45 Move to Pb-D1 to finish trigger study with narrow window.
- 05:53 Start run 1723. Window study of new, narrower trigger
- 06:06 Start run 1724. Mostly coincidences for window study.
- 06:15 Start run 1725, 0.5 uA rate study. Mostly T3.
- 06:26 Start run 1726, 1.0 uA rate study. Mostly T3.
- 06:33 Start run 1727, 2.0 uA rate study. Mostly T3.
- 06:45 MCC takes beam to study BLMs for a few minutes after run
ends. South Linac has an unstable RF zone.
- 07:30 MCC calls with an update. The south linac is having a cryo
problem and they have called in techs. Beam will be a while.